A 0.05 significance level is used for a hypothesis test of the claim that when parents use a particular method of gender​ selection, the proportion of baby girls is less than 0.5. Assume that sample data consists of 78 girls in ​169 births, so the sample statistic of 6/13 results in a z score that is 1 standard deviation below 0. Complete parts​ (a) through​ (h) below.

a. Identify the null hypothesis and the alternative hypothesis:

b. What is the value of a?

a=___

c. What is the sampling distribution of the sample statistic?

__Normal distribution

__x^2

__Student (t) distribution

d. is the test two-tailed, left-tailed, or right-tailed?

e. what is the value of the test statistic?

f. What is the P-value? (Round four decimals)

g. What are the critical value/s? (Round two decimals)

h. What is the area of the critical region? (Round two decimals)
Answer #1

Given : n=169 , x=78

The estimate of the sample proportion is , \hat{p}=\frac{x}{n}=\frac{78}{169}=0.4615

a. The null and alternative hypothesis is ,

H_0:p=0.5;H_1:p<0.5

b. The value of a is a=0.05

c. The sampling distribution of the sample statistic is Normal distribution

d. The test is left-tailed test

e. The value of the test statistic is ,

Z_{stat}=\frac{\hat{p}-p_0}{\sqrt{\frac{p_0(1-p_0)}{n}}}=\frac{0.4615-0.5}{\sqrt{\frac{0.5(1-0.5)}{169}}}=-1

f. The p-value is ,

p-value=P(Z>|Z_{stat}|)=P(Z>1)=1-P(Z\leq 1)=1-\Phi(1)

=1-0.8413=0.1587 ; From standard normal distribution table

g. The critical value is ,

Za 20.05 -1.64 ; From Z-table

h. The area of the critical region is (-\infty ,-1.64]